OpenStudy (anonymous):

Write the following equation in the general form Ax + By + C = 0. 2x + y = 6 2x + y - 6 = 0 2x + y + 6 = 0 2x - y - 6 = 0

OpenStudy (anonymous):

Um... Except for the first one, they're all written in general form

OpenStudy (aravindg):

ya

OpenStudy (aravindg):

take 6 to left

OpenStudy (anonymous):

You're starting with 2x+y=6 and trying to rewrite it in general form. That means you need to get the 6 to the other side of the equals sign. How can you move a constant from one side of an equation to another?

OpenStudy (anonymous):

Oh, that first one is the problem. Silly me!!!

OpenStudy (anonymous):

2x - 3y - 6 = 0

OpenStudy (anonymous):

\[2x - 3y - 6 = 0\]

OpenStudy (aravindg):

\[\huge 2x+y-6\]

OpenStudy (anonymous):

?

OpenStudy (aravindg):

2x+y-6=0

OpenStudy (anonymous):

@AravindG how do u increace the font size! plz teach

OpenStudy (aravindg):

\[\huge 2x+y-6=0\]

OpenStudy (aravindg):

@Eric_Playsinacoolway its easy use \huge

OpenStudy (anonymous):

\huge 2x+y-6=0

OpenStudy (anonymous):

Well that didn't work.

OpenStudy (anonymous):

\[\huge2x - 3y - 6 = 0\]

OpenStudy (aravindg):

hey its not -3

OpenStudy (aravindg):

@nbouscal its understood

OpenStudy (anonymous):

\huge thanks @AravindG and @nbouscal .

OpenStudy (anonymous):

Still not working...:(

OpenStudy (anonymous):

\[\huge I think I got \it. :)\]

OpenStudy (anonymous):

*it.